# Object Segmentation-Assisted Inter Prediction for Versatile Video Coding

### v1

@@ VII-C Performance Analysis
| Dataset | Sequence Num | Objects Num1 | Selected Character | LDB (R2 /L3 ) | RA (R/L) | Accuracy (R-L)4 | LDB 6 | RA 6 | ||||
| Average5 | Highest5 | Average | Highest | & Mean4 | Average | Highest | Average | Highest | ||||
| DAVIS-17[48] | 30 | 99 | Diverse Scenes Multiple Objects Complex Motions High Quality | -0.21%/-0.51% | -0.36%/-0.81% | -0.11%/-0.39% | -0.33%/-0.72% | 83.2/83.1/81.6/77.7 | -0.30% | -0.45% | -0.28% | -0.39% |
| Youtube-VOS[49] | 30 | 120 | -0.19%/-0.41% | -0.31%/-0.77% | -0.04%/-0.31% | -0.22%/-0.64% | 80.7/79.3/77.3/74.9 | -0.22% | -0.46% | -0.27% | -0.42% | |
| MOSE[50] | 30 | 295 | -0.16%/-0.43% | -0.41%/-0.73% | 0.09%/-0.27% | -0.25%/-0.53% | 72.3/68.5/62.7/57.2 | -0.27% | -0.32% | -0.36% | -0.28% | |
| Overall | 90 | 514 | -0.19%/-0.45% | -0.36%/-0.77% | -0.02%/-0.32% | -0.27%/-0.63% | 78.7/77.0/73.9/69.9 | -0.26% | -0.41% | -0.30% | -0.36% | |
1 Objects Num denotes the total object number of the test sequences. The object number of each test sequence is shown in Section V.B of Supplementary Material.
2 R indicates the test results are performed by the proposed reference frame-based segmentation method (Section III of supplementary material).
@@
VII-C1 Ablation Study
First, we validate the effectiveness of SA-MC and study the influence of different components of SA-MC, including the two-step prediction fusion (TWO-STEP) and overlapped region-based weighted strategy (ORMC). As shown in Table V , we introduce two variants for comparison with SA-MC. ONE-STEP indicates that the final prediction is directly derived according to the segmentation mask from the block-level primary and secondary predictions at the end of fractional-part MC. TWO-STEP indicates that prediction fusion is performed on the integer and fractional parts of MC. The integer-part fusion mainly aims to generate the accurate integer pixels for the subsequent fractional MC process, and provides the precise adjacent reference pixels for the interpolation of edge-adjacent fractional pixels. SAIP indicates that the TWO-STEP and ORMC are all used in SA-MC. Through the experimental results, we draw two conclusions: (1) For the comparison of ONE-STEP and SAIP, the experimental results show that SA-MC is superior to the direct multi-region MC, which demonstrates the assistance of a segmentation mask is beneficial to generating accurate prediction. (2) For the comparison of two variants and SAIP, the experimental results show that TWO-STEP and ORMC have a significant effect on SA-MC, and further demonstrate that TWO-STEP can better assist DCTIF in generating accurate fractional-pixel prediction and ORMC is beneficial to reducing the boundary artifacts.
Second, we validate the effectiveness of SA-MVC. In Table V , we introduce a variant for comparison with SA-MVC. Instead of using SA-MVC, the secondary candidate list is constructed according to the regular merge candidate order. For the comparison of variant and SAIP, the experimental results show that our proposed MVC method helps achieve more bits saving than the regular method, which demonstrates the assistance of a segmentation mask is beneficial to coding the MV efficiently.
VII-C2 Comparison of Mode Selected Blocks
Figure 7 shows the mode selected results of the GEO mode and our proposed SAIP mode, more results are shown in Section VI of supplementary material. Compared with GEO, SAIP possesses more flexible motion field description capability, which focuses on more motion-inconsistent regions at the edges of moving objects and handles more difficult cases such as the outer outline of the people in BQMall. Meanwhile, the SAIP not only tackles the efficient representation of complex motion fields, but also tends to hit the larger block to save the bit consumption of the refined partition.
VII-D Limitation and Potential Analysis
VII-D1 Limitation
@@
Based on the above experiments, we find that the performance of SAIP is limited by segmentation accuracy. Here we analyze it in two aspects: (1) Segmentation Technology: From the CTC results of SAIP (Table I ), we find that our method performs poorly on some sequences with special cases (extreme motion, serious occlusion, large resolution), which is mainly limited by the poor segmentation results, such as BQTerrence (camera motion), Tango2 (indistinguishable foreground and background), Campfire (dense moving objects). (These failure segmentation cases are shown in Section V.A of supplementary material.) Meanwhile, learned segmentation schemes have not yet been widely generalized to large-resolution video, such as 1080P, 4k (resolution of Class A, B). The datasets and benchmarks [48, 49] of these segmentation schemes are mainly evaluated on medium resolution (480P-720P). Due to the limited capability and generalization of the learned segmentation method [27, 33], it’s difficult to generate accurate segmentation to assist the motion prediction of SAIP in some challenging scenes. (2) Dependence of Temporal-domain Reference Structure: From the comparison of SAIP’s CTC results under different configurations in Table I , we find that the reference structure also affects its performance. Due to the motion area partition of each block being derived from the reference block, the motion area partition of the long-distance reference block is difficult to adapt to the motion situation of the current coding block and further influences prediction accuracy, such as the RA reference structure case.
VII-D2 Potential
@@
To explore the potential improvement brought by segmentation accuracy, we test the SAIP on the datasets (train and validation datasets of VOS Benchmark [48, 49, 50]) with hand-annotated label segmentation mask to avoid the above limitations. For each dataset, we select 30 videos with multiple moving objects, and test the SAIP under different segmentation settings, including the proposed segmentation method (Section III of supplementary material) and hand-annotated label segmentation mask. The detailed settings and experimental results are shown in Table VII . Table VII shows the difference of SAIP’s performance under different segmentation settings (LDB/RA ) can achieve up to -0.41%, -0.36%, and on average -0.26%, -0.30% BD-rate reduction (Y component) on VTM-12.0. (The visualization of the different settings’ segmentation results is shown in Section V.B of supplementary material.) It demonstrates that the SAIP can achieve higher performance with the assistance of the better segmentation mask, and exhibits positive performance potential with continuous enhancement of advanced segmentation technology in the future.
VIII Conclusion

### v2

@@ IX-C Performance Analysis
IX-C1 Ablation Study
First, we validate the effectiveness of SA-MC and study the influence of different components of SA-MC, including the two-step prediction fusion (TWO-STEP) and overlapped region-based weighted strategy (ORMC). As shown in Table VIII , we introduce two variants for comparison with SA-MC. ONE-STEP indicates that the final prediction is directly derived according to the segmentation mask from the block-level primary and secondary predictions at the end of fractional-part MC. TWO-STEP indicates that prediction fusion is performed on the integer and fractional parts of MC. The integer-part fusion mainly aims to generate the accurate integer pixels for the subsequent fractional MC process, and provides the precise adjacent reference pixels for the interpolation of edge-adjacent fractional pixels. SAIP indicates that the TWO-STEP and ORMC are all used in SA-MC. Through the experimental results, we draw two conclusions: (1) For the comparison of ONE-STEP and SAIP, the experimental results show that SA-MC is superior to the direct multi-region MC, which demonstrates the assistance of a segmentation mask is beneficial to generating accurate prediction. (2) For the comparison of two variants and SAIP, the experimental results show that TWO-STEP and ORMC have a significant effect on SA-MC, and further demonstrate that TWO-STEP can better assist DCTIF in generating accurate fractional-pixel prediction and ORMC is beneficial to reducing the boundary artifacts.
Second, we validate the effectiveness of SA-MVC. In Table VIII , we introduce a variant for comparison with SA-MVC. Instead of using SA-MVC, the secondary candidate list is constructed according to the regular merge candidate order. For the comparison of variant and SAIP, the experimental results show that our proposed MVC method helps achieve more bits saving than the regular method, which demonstrates the assistance of a segmentation mask is beneficial to coding the MV efficiently.
IX-C2 Comparison of Mode Selection
Figure 9 shows the mode selected results of the GEO mode and our proposed SAIP mode, more results are shown in Section III of supplementary material. Compared with GEO, SAIP possesses more flexible motion field description capability, which focuses on more motion-inconsistent regions at the edges of moving objects and handles more difficult cases such as the outer outline of the people in BQMall. Meanwhile, the SAIP not only tackles the efficient representation of complex motion fields, but also tends to hit the larger block to save the bit consumption of the refined partition.
IX-D Limitation and Potential
| Dataset | Sequence Num | Objects Num1 | Selected Character | LDB (R2 /L3 ) | RA (R/L) | Accuracy (R-L)4 | LDB 6 | RA 6 | ||||||||||||
| Average5 | Highest5 | Average | Highest | & Mean4 | Average | Highest | Average | Highest | ||||||||||||
| DAVIS-17[52] | 30 | 99 | Diverse Scenes Multiple Objects Complex Motions High Quality | -0.21%/-0.51% | -0.36%/-0.81% | -0.11%/-0.39% | -0.33%/-0.72% | 83.2/83.1/81.6/77.7 | -0.30% | -0.45% | -0.28% | -0.39% | ||||||||
| Youtube-VOS[53] | 30 | 120 | -0.19%/-0.41% | -0.31%/-0.77% | -0.04%/-0.31% | -0.22%/-0.64% | 80.7/79.3/77.3/74.9 | -0.22% | -0.46% | -0.27% | -0.42% | |||||||||
| MOSE[54] | 30 | 295 | -0.16%/-0.43% | -0.41%/-0.73% | 0.09%/-0.27% | -0.25%/-0.53% | 72.3/68.5/62.7/57.2 | -0.27% | -0.32% | -0.36% | -0.28% | |||||||||
| Overall | 90 | 514 | -0.19%/-0.45% | -0.36%/-0.77% | -0.02%/-0.32% | -0.27%/-0.63% | 78.7/77.0/73.9/69.9 | -0.26% | -0.41% | -0.30% | -0.36% | |||||||||
1 Objects Num denotes the total object number of the test sequences. The object number of each test sequence is shown in Section II-B of Supplementary Material.
2 R indicates the test results are performed by the proposed reference frame-based segmentation method (Section IV).
@@ IX-D Limitation and Potential
IX-D1 Limitation
@@
Based on the above experiments, we find that the performance of SAIP is limited by segmentation accuracy. Here we analyze it in two aspects: (1) Segmentation Technology: From the CTC results of SAIP (Table II ), we find that our method performs poorly on some sequences with special cases (extreme motion, serious occlusion, large resolution), which is mainly limited by the poor segmentation results, such as BQTerrence (camera motion), Tango2 (indistinguishable foreground and background), Campfire (dense moving objects). (These failure segmentation cases are shown in Section II-A of supplementary material.) Meanwhile, learned segmentation schemes have not yet been widely generalized to large-resolution video, such as 1080P, 4k (resolution of Class A, B). The datasets and benchmarks [52, 53] of these segmentation schemes are mainly evaluated on medium resolution (480P-720P). Due to the limited capability and generalization of the learned segmentation method [31, 37], it’s difficult to generate accurate segmentation to assist the motion prediction of SAIP in some challenging scenes. (2) Dependence of Temporal-domain Reference Structure: From the comparison of SAIP’s CTC results under different configurations in Table II , we find that the reference structure also affects its performance. Due to the motion area partition of each block being derived from the reference block, the motion area partition of the long-distance reference block is difficult to adapt to the motion situation of the current coding block and further influences prediction accuracy, such as the RA reference structure case.
IX-D2 Potential
@@
To explore the potential improvement brought by segmentation accuracy, we test the SAIP on the datasets (train and validation datasets of VOS Benchmark [52, 53, 54]) with hand-annotated label segmentation mask to avoid the above limitations. For each dataset, we select 30 videos with multiple moving objects, and test the SAIP under different segmentation settings, including the proposed segmentation method (Section IV) and hand-annotated label segmentation mask. The detailed settings and experimental results are shown in Table IX . Table IX shows the difference of SAIP’s performance under different segmentation settings (LDB/RA ) can achieve up to -0.41%, -0.36%, and on average -0.26%, -0.30% BD-rate reduction (Y component) on VTM-12.0. (The visualization of the different settings’ segmentation results is shown in Section II-B of supplementary material.) It demonstrates that the SAIP can achieve higher performance with the assistance of the better segmentation mask, and exhibits positive performance potential with continuous enhancement of advanced segmentation technology [16, 17, 18] in the future.
X Conclusion